Instructions

Please fill the user’s information on the request form, and sign it at the bottom. The user’s supervisor must fill the supervisor’s information and sign the form. In case the user holds a permanent position within the Institution specified on the form, he/she may fill and sign as his/her own supervisor.

Please return the completed form by post to

Alfredo Ferrari EN/STI, Mailbox Z04500 CERN CH-1211 Geneva 23 Switzerland fax: +41-227669474

Sending of a copy by fax may speed up the procedure, however the original form must also be sent by standard mail.

Request form for the FLUKA code

March 3, 2011

Single-User , Single-site information

Name  HU Liqin
FLUKA user id fuid-3308

Institution

Address

e-mail  lhu@ipp.ac.cn

Liason Officer or Supervisor

Name

e-mail

By signature below, the subscriber acknowledges and agrees to comply with the FLUKA User license and the associated conditions as reported below.

The subscriber acknowledges and agrees that the FLUKA software is licensed, on a non-exclusive basis, for use only by the identified single user only on computers owned or leased by the user’s Institution as designated above. The license expires automatically when the user leaves his/her position in the specified Institution. Access by anyone other than the identified user is strictly prohibited. The subscriber declares and agrees that the FLUKA code will not be used directly or indirectly in any or more of the following areas

• Weapons or their precursors, and associated hardware, software or technical data

• Commercial activities

• Training, unless fee-less and restricted to academic/educational audience.

The licensee declares that he/she has carefully read the release notes and he/she accepts the associated recommendations. In particular the licensee acknowledges that

• The availability of the source code shall not be exploited for tentative builds on other architectures or with different compilers/compiler options than the ones recommended by the development team.

• The licensee shall regularly update his FLUKA version according to the FLUKA release notes. The availability of the source code shall not be exploited to circumvent the expiration date.

FLUKA User license

COPYRIGHT NOTICE AND LICENSE CONDITIONS

Copyright Italian National Institute for Nuclear Physics (INFN) and European Organization for Nuclear Research (CERN), 1989–2010.

All rights not expressly granted under this license are reserved.

Requests for permissions not granted under this license shall be addressed to the FLUKA Collaboration Committee, through fcc@fluka.org. Any permission may only be granted in writing.

This software results in particular from work performed by Alberto Fassò, Alfredo Ferrari, Johannes Ranft, Paola Sala (the “Authors”), and their collaborators (the “Collaborators”).

INFN and CERN are the exclusive source of distribution of the code, bug fixes and documentation of the FLUKA software (http://www.fluka.org), and may authorise distribution by mirror sites.

This license cancels and replaces any prior license conditions but their warranty and liability provisions shall continue to apply to any use or modifications made under such prior license conditions.

DEFINITIONS

The FLUKA software (“FLUKA”) means the fully integrated particle physics Monte Carlo simulation software package being developed since 1989, available from the official FLUKA website http://www.fluka.org and authorised mirror sites. FLUKA is made up of FLUKA core code and FLUKA User Routines.

The FLUKA Copyright Holders means both CERN and INFN.

FLUKA User Routines means the set of subroutines collected in the usermvax section of FLUKA and forming part of the standard distribution of FLUKA.

The Licensee means any person acting individually within a non-profit organisation, exercising any permission granted by this license.

LICENSE GRANT

1. Subject to the terms and conditions of this license, the FLUKA Copyright Holders herewith grant to the Licensee a worldwide, non-exclusive, royalty-free, source and object code license to use and reproduce FLUKA for internal scientific non commercial non-military purposes only. Notwithstanding the foregoing, the Licensee shall not execute FLUKA in a manner that produces an output whose contents are directly useable or easily employable to simulate the physics models embedded within FLUKA in a generic manner, or excise portions of FLUKA source or object code, and execute them independently of FLUKA. Extracting specific isolated results from any of the individual internal physics models embedded within FLUKA is not permitted. Permitted use and reproduction are referred to below as “Use”.

2. Modification (including translation) of FLUKA, in whole or in part, is not permitted, except for modification of FLUKA User Routines that do not circumvent, replace, add to or modify any of the functions of the FLUKA core code. Permitted modifications are referred to below as ”Modifications”.

3. FLUKA is licensed for Use by the Licensee only, and the Licensee shall not market, distribute, transfer, license or sub-license, or in any way make available (“Make Available”) FLUKA or Modifications, in whole or in part, to third parties, without prior written permission. The Licensee shall not assign or transfer this license.

4. Notwithstanding section 3, the Licensee may Make Available his Modifications of FLUKA User Routines to third parties under these license conditions.

5. The Licensee shall not insert FLUKA code or Modifications, in whole or in part, into other codes without prior written permission.

6. Any use of FLUKA outside the scope of this license is subject to prior written permission.

GRANT BACK

7. The Licensee shall in a timely fashion notify to fcc@fluka.org any Modifications carried out by him. Except for Authors, Collaborators, and employees of the FLUKA Copyright Holders, the copyright in whose Modifications shall automatically be vested in the FLUKA Copyright Holders, the Licensee herewith grants the FLUKA Copyright Holders a perpetual, royalty- free, irrevocable and non-exclusive license to his Modifications, with no limitation of use. The Licensee acknowledges that the FLUKA Copyright Holders may insert such Modifications into future releases of FLUKA, subject to appropriate acknowledgment of the Licensee’s contribution.

8. The Licensee shall report as soon as practical any errors or bugs found in any portion

of FLUKA to fluka-discuss@fluka.org

PUBLICATIONS AND ACKNOWLEDGEMENT

9. The Licensee shall explicitly acknowledge his use of FLUKA in any publication or communication, scientific or otherwise, relating to such use, by citing the FLUKA set of references (http://www.fluka.org^{1}) and the FLUKA copyright notice.

1

10. The Licensee shall ensure that the FLUKA set of references, the FLUKA copyright notice and these license conditions are not altered or removed from FLUKA and that all embodiments of FLUKA and Modifications contain in full the FLUKA set of references, the FLUKA copyright notice, and these license conditions.

11. Any insertion of FLUKA code or Modifications, in whole or in part, into other codes with permission under section 5 shall preserve the FLUKA set of references, the FLUKA copyright notice and these license conditions in the FLUKA code or Modifications concerned, and must also reproduce these within any additional global notices included along or embedded within the software into which the FLUKA code or the Modifications have been integrated, in whole or in part. Any part of the FLUKA code or Modifications so inserted shall continue to be subject to these license conditions.

12. Publication of any results of comparisons of specific internal physics models extracted from FLUKA with permission under section 6 with data or with other codes or models is subject to prior written permission.

13. Contributions to any formal code comparisons and validation exercises pertaining to FLUKA, sponsored by recognised bodies or within the framework of recognised conferences and workshops, are subject to prior written permission.

WARRANTY AND LIABILITY

14. DISCLAIMER FLUKA IS PROVIDED BY THE FLUKA COPYRIGHT HOLD-ERS “AS IS” AND ANY EXPRESS OR IMPLIED WARRANTIES, INCLUDING, BUT NOT LIMITED TO, IMPLIED WARRANTIES OF MERCHANTABILITY, OF SATISFACTORY QUALITY, AND FITNESS FOR A PARTICULAR PUR-POSE OR USE ARE DISCLAIMED. THE FLUKA COPYRIGHT HOLDERS AND THE AUTHORS MAKE NO REPRESENTATION THAT FLUKA AND MODIFICATIONS THEREOF WILL NOT INFRINGE ANY PATENT, COPY-RIGHT, TRADE SECRET OR OTHER PROPRIETARY RIGHT.

• A. Ferrari, P.R. Sala, A. Fassò, and J. Ranft, “FLUKA: a multi-particle transport code”, CERN-

2005-10 (2005), INFN/TC 05/11, SLAC-R-773.

• G. Battistoni, S. Muraro, P.R. Sala, F. Cerutti, A. Ferrari, S. Roesler, A. Fassò, J. Ranft, “The

FLUKA code: Description and benchmarking”, Proceedings of the Hadronic Shower Simulation Workshop 2006, Fermilab 6–8 September 2006, M. Albrow, R. Raja eds., AIP Conference Proceeding 896, 31-49, (2007)

15. LIMITATION OF LIABILITY THE FLUKA COPYRIGHT HOLDERS AND ANY CONTRIBUTOR SHALL HAVE NO LIABILITY FOR DIRECT, INDIRECT, SPECIAL, INCIDENTAL, CONSEQUENTIAL, EXEMPLARY, PUNITIVE OR OTHER DAMAGES OF ANY CHARACTER INCLUDING, WITHOUT LIMI-TATION, PROCUREMENT OF SUBSTITUTE GOODS OR SERVICES, LOSS OF USE, DATA OR PROFITS, OR BUSINESS INTERRUPTION, HOWEVER CAUSED AND ON ANY THEORY OF CONTRACT, WARRANTY, TORT (IN-CLUDING NEGLIGENCE), PRODUCT LIABILITY OR OTHERWISE, ARIS-ING IN ANY WAY OUT OF THE USE OF FLUKA, EVEN IF ADVISED OF THE POSSIBILITY OF SUCH DAMAGES, AND THE LICENSEE SHALL HOLD THE COPYRIGHT HOLDERS AND ANY CONTRIBUTOR FREE AND HARM-LESS FROM ANY LIABILITY, INCLUDING CLAIMS BY THIRD PARTIES, IN RELATION TO SUCH USE.

TERMINATION

16. This license shall terminate with immediate effect and without notice if the Licensee fails to comply with any of the terms of this license, or if the Licensee initiates litigation against any of the FLUKA Copyright Holders or any contributors with regard to FLUKA. It shall also terminate with immediate effect from the date on which a new version of FLUKA becomes available. In either case sections 14 and 15 above shall continue to apply to any Use or Modifications made under these license conditions.

The use of the FLUKA code must be acknowledged explicitly by quoting the following references:

• A. Ferrari, P.R. Sala, A. Fassò, and J. Ranft, “FLUKA: a multi-particle transport code”, CERN-2005-10 (2005), INFN/TC 05/11, SLAC-R-773.

• G. Battistoni, S. Muraro, P.R. Sala, F. Cerutti, A. Ferrari, S. Roesler, A. Fassò, J. Ranft, “The FLUKA code: Description and benchmarking”, Proceedings of the Hadronic Shower Simulation Workshop 2006, Fermilab 6–8 September 2006, M. Albrow, R. Raja eds., AIP Conference Proceeding 896, 31-49, (2007)

Additional FLUKA references can be added, provided they are relevant for the FLUKA version under consideration.

This set of references is subject to change in time. New ones will be communicated, when necessary, in the Release Notes of new FLUKA versions. Note:

The Nuclear Energy Agency Data Bank is an authorized distributor of the code

For questions/clarifications/problems about the license or in general the use of FLUKA please contact the joint INFN/CERN FLUKA Collaboration Committee at fcc@fluka.org, or one of the Authors.

Signature:

print name  signature  date

and as supervisor:

print name  signature  date